A proposed steel bridge has an indefinite life. The initial cost of the bridge is $4,750,000 and annual maintenance costs are estimated to be $25,000. The bridge deck will be resurfaced every 10 years for $900,000, and anti-corrosion paint will be applied every 5 years for $250,000. If the interest rate is 8 % what is the EAC? If 650,000 axles will cross the bridge each year what approximate toll should be charged to the nearest nickel (round up to nearest nickel…you would want cover all your costs).
Answer #1

EAC = 4,750,000(A/P, 8%, infinity) + 900,000(A/F, 8%, 10) + 250,000(A/F, 8%, 5) + 25,000

        = 4,750,000(0.08) + 900,000(0.0690) + 250,000(0.1705) + 25,000

       = 380,000 + 62,000 + 25,000

       = $467,100

Toll = 467,100 / 650,000

      = $0.72 per axles